# SpikingBrain: Spiking Brain-inspired Large Models

### v2

@@ 4.1 Distributed Training Adaptation
  • Multi-Granularity Recomputation (korthikanti2022selectiverecompute, ): To balance computation and memory under high memory pressure, experts support three recomputation levels: lightweight (activations and router), moderate (including FFN and shared experts), and full (entire MoE layer).
@@
To address high communication overhead in distributed training, MetaX utilizes SDMA engines for intra-node high-speed data transfer. For tensor parallelism and expert parallelism, communication kernels are fused with compute kernels to reduce scheduling conflicts (NVIDIA2025overlap, ). Memory optimizations include fine-grained offloading of transformer layer weights or optimizer states to the CPU, as well as selective recomputation of normalization layers, activation functions, or individual transformer layers based on memory demand.
@@
  • Fast Checkpointing: The DLRover (wang2023dlrover, ) Flash Checkpoint technique first writes training state (model weights, optimizer, learning rate scheduler, etc.) to CPU memory before asynchronously persisting to a distributed file system. This reduces the I/O time by 85% and shortens recovery time after failures. The built-in profiling tools automatically instrument training jobs, monitor performance per layer, detect slow nodes, and trigger alerts, helping maintain high cluster utilization.
@@ 4.2 Operator Adaptation
The efficient adaptation of SpikingBrain large models to MetaX GPUs relies on the comprehensive MetaX software ecosystem. The overall process can be divided into two major components: Triton adaptation and CUDA migration to the MetaX self-developed MACA framework. These two pathways are designed for different subsets of operators within SpikingBrain models. While distinct in implementation, they complement each other and together constitute a complete hardware adaptation framework tailored for MetaX GPUs (see Figure 5).
@@
  • JIT Compilation Optimization: During Triton’s just-in-time compilation stage, the linear attention operators (yang2024gated, ) used in SpikingBrain are reorganized at the code level. By refining instruction pipelining and register allocation, the operators achieve a balance between memory access latency and computational density. This stage directly leverages Triton’s hardware-agnostic general optimization strategies, enabling dynamic and efficient execution of compute kernels.
@@
  • Target Code Generation via MetaX Compiler: In the final stage, Triton kernels are transformed by the MetaX compiler backend into executable target machine code for MetaX GPUs. Beyond standard code generation, the compiler introduces deep optimizations for tensor cores, SIMD instruction sets, and memory alignment constraints, ensuring operators run at near-peak hardware performance.
@@
It is worth emphasizing that the entire adaptation and migration process for brain-inspired large models on MetaX GPUs adheres to a user-friendly design philosophy. Whether through Triton kernel optimization or CUDA-to-MACA migration, end users can preserve their existing programming practices and interface calls. In most cases, no extensive code modifications are required for the models to run efficiently on MetaX GPUs. Meanwhile, the MetaX ecosystem provides unified debugging and performance analysis tools, enabling developers to transparently monitor execution characteristics on hardware and further fine-tune performance as needed.
Figure 5: Operator adaptation of SpikingBrain on MetaX GPUs. The adaptation involves two complementary pathways: Triton adaptation and CUDA migration to the MACA framework, covering different operator subsets. Together, they form a unified hardware adaptation framework tailored for MetaX GPUs.
@@ 4.3 Parallel Topology
Data parallelism (Valiant1990Bridging, ) involves partitioning the training data into batches, each processed by a separate GPU. Each GPU maintains a complete replica of the model and performs forward and backward passes independently, while gradient synchronization occurs only at the beginning of the backward propagation. Data parallelism offers low communication overhead and excellent scalability. As computing resources increase, the scale of data parallelism can grow accordingly, increasing the total batch size to enhance overall throughput and reduce training time. During training, we employ ZeRO (rajbhandari2020zero, ) to eliminate redundancy of the optimizer states, effectively distributing the GPU memory pressure.
@@
Pipeline parallelism divides the model into different stages by layers. Each GPU is assigned a subset of layers, and intermediate activations are communicated between GPUs during the forward and backward passes using peer-to-peer (p2p) communication. We employ the 1F1B (narayanan2021efficientllmmg, ; harlap2018pipedream, ) scheduling algorithm for efficient pipeline execution. Furthermore, incorporating Mixture-of-Experts (MoE) layers interspersed with dense FFN layers in SpikingBrain-76B helps mitigate the memory imbalance issue, where the first pipeline stage typically exhibits higher memory usage compared to other stages.
@@
In Mixture-of-Experts models, most parameters reside within the expert networks. Expert parallelism (rajbhandari2022deepspeed, ) addresses this by partitioning experts across multiple GPUs. This requires all-to-all communication to dispatch and combine tokens. During the training of SpikingBrain-76B, we employ Grouped GEMM kernels to accelerate computation across experts, and an auxiliary loss is used to promote balanced token routing across experts.
@@
Training with long sequences incurs substantial memory costs for storing intermediate activations. Sequence parallelism tackles this by splitting the input sequence along its length dimension, with each GPU processes a contiguous chunk. Communication primitives like all-to-all (a2a) or peer-to-peer (p2p) are then used to exchange necessary information between devices, enabling the computation of attention over the entire sequence. For our experiments, we employ DeepSpeed Ulysses (jacobs2023deepspeed, ) for efficient sequence parallelism and incorporate specialized algorithms for linear attention modules to further improve performance. For linear attention, AllGather communication is used to collect intermediate local states (Sun2025LASP2RS, ) when the SP size is small, while for larger SP sizes (e.g., across multiple nodes), the All-Scan primitive from ZeCO (chou2025zeco, ) sequence parallelism is adopted to effectively eliminate communication overhead.
@@
The SpikingBrain-7B model is trained using the Colossal-AI (li2023colossal, ) framework. For 128k sequence length training, a parallel strategy of 32-way data parallelism and 8-way sequence parallelism is adopted. To reduce GPU memory consumption, techniques such as ZeRO-2 (rajbhandari2020zero, ) and activation recomputation (chen2016recompute, ; korthikanti2022selectiverecompute, ) are applied. Sequence parallelism is implemented with DeepSpeed Ulysses (jacobs2023deepspeed, ), which performs all-to-all communication within nodes. In addition, the number of attention heads is padded to satisfy the partition requirements.
The SpikingBrain-76B model is trained using the Megatron (shoeybi2020megatronlm, ) framework. For 8k sequence length training, the parallel configuration consists of 128-way data parallelism, 8-way expert parallelism, and 4-way pipeline parallelism. ZeRO (rajbhandari2020zero, ) and selective activation recomputation (chen2016recompute, ; korthikanti2022selectiverecompute, ) are applied to reduce memory consumption. For extended contexts of 32k and 128k, 4-way and 8-way sequence parallelism are added on top of the base 8k parallel strategy, respectively. The 76B hybrid model integrates linear attention and (local) softmax attention. The softmax attention branch uses DeepSpeed Ulysses (jacobs2023deepspeed, ) with all-to-all communication and head padding to enable partitioning, while the linear attention branch employs the AllGather-based SP algorithm (LASP-2 (Sun2025LASP2RS, )) to reduce communication overhead.
@@ 5.1 Downstream Performance
We conduct a comprehensive evaluation of the two converted efficient models, SpikingBrain-7B and -76B, across a broad set of downstream tasks [^3] This section evaluates the effectiveness of the architectural design and conversion pipeline. All models remain in floating-point format; results for spiking models are presented in Section 5.5.. They are compared against open-source baselines of similar scale, including Hybrid, Transformer, and MoE architectures, as well as the strong base model Qwen2.5-7B. To ensure fairness, all evaluations are performed under identical settings using the OpenCompass (buitrago2019open, ) framework. In selecting evaluation metrics, we place greater emphasis on pretraining-oriented general-purpose benchmarks (e.g., MMLU, CMMLU, C-Eval, ARC-C, HS), as these better indicate whether our models—trained with fewer than 200B tokens during efficient conversion—can inherit the knowledge and modeling capacity of the base model. Moreover, these benchmarks reveal whether the converted models preserve strong generalization while demonstrating significant long-sequence efficiency advantages, without being confounded by data quality factors such as alignment performance after SFT.
As shown in Table 1, our SpikingBrain-7B linear model recovers nearly 90% of the base model’s performance across benchmarks, reaching a level comparable to advanced Transformer models such as Mistral-7B and Llama3-8B. This indicates that, with appropriate training strategies, linear attention architectures can preserve strong general-purpose modeling ability while substantially reducing inference complexity. However, due to the significant structural modifications of the attention mechanism, there remains a certain performance gap between the 7B pure linear model and base Qwen2.5-7B. This suggests that the extent of recovery is strongly influenced by architectural changes when pursuing extreme efficiency.
Table 1: Performance evaluation of the SpikingBrain-7B pre-trained model. All models are tested with the HuggingFace framework and evaluated using a perplexity-based method. Except for Qwen2.5, the other baselines are trained on limited Chinese data, resulting in clear disadvantages on CMMLU and C-Eval.
@@
MMLU (hendrycks2020mmlu, ) 65.84 63.24 62.56 58.19 65.74 74.21
CMMLU (li2023cmmlu, ) 71.58 42.50 44.58 38.42 52.44 81.73
ARC-C (clark2018arcc, ) 43.32 47.53 45.13 37.18 51.96 44.04
HS (zellers2019hellaswag, ) 70.89 71.50 75.81 52.02 71.60 72.81
Ceval (huang2023ceval, ) 69.80 41.93 47.04 36.40 51.46 81.60
@@
Table 2: Performance evaluation of the SpikingBrain-76B pre-trained model. All models are tested with the vLLM (kwon2023vllm, ) framework and evaluated using a perplexity-based method. Except for Qwen2.5, the other baselines are trained on limited Chinese data, resulting in clear disadvantages on CMMLU and C-Eval.
@@
SpikingBrain-76B
(76B-A12B)
[…]
Jamba (lieber2024jamba, )
(52B-A12B)
[…]
Mixtral-8*7B (jiang2024mixtral, )
(47B-A13B)
[…]
Qwen2.5 (Yang2024Qwen2TR, )
@@
MMLU (hendrycks2020mmlu, ) 73.58 67.17 71.23 69.57 75.94 75.31
CMMLU (li2023cmmlu, ) 78.83 51.11 52.70 52.94 61.80 81.50
ARC-C (clark2018arcc, ) 42.00 49.10 48.98 46.21 57.64 43.56
HS (zellers2019hellaswag, ) 73.31 79.39 79.42 79.15 79.34 73.37
Ceval (huang2023ceval, ) 78.89 48.94 54.39 49.26 61.02 81.68

### v3

@@ 4.1 Distributed Training Adaptation
  • Multi-Granularity Recomputation (Korthikanti et al., 2022): To balance computation and memory under high memory pressure, experts support three recomputation levels: lightweight (activations and router), moderate (including FFN and shared experts), and full (entire MoE layer).
@@
To address high communication overhead in distributed training, MetaX utilizes SDMA engines for intra-node high-speed data transfer. For tensor parallelism and expert parallelism, communication kernels are fused with compute kernels to reduce scheduling conflicts (NVIDIA, 2025). Memory optimizations include fine-grained offloading of transformer layer weights or optimizer states to the CPU, as well as selective recomputation of normalization layers, activation functions, or individual transformer layers based on memory demand.
@@
  • Fast Checkpointing: The DLRover (Wang et al., 2023) Flash Checkpoint technique first writes training state (model weights, optimizer, learning rate scheduler, etc.) to CPU memory before asynchronously persisting to a distributed file system. This reduces I/O time by 85% and shortens recovery time after failures. The built-in profiling tools automatically instrument training jobs, monitor performance per layer, detect slow nodes, and trigger alerts to sustain high cluster utilization.
Figure 5: Operator adaptation of SpikingBrain on MetaX GPUs. The adaptation involves two complementary pathways: Triton adaptation and CUDA migration to MACA framework, covering different operator subsets. Together, they form a unified hardware adaptation framework tailored for MetaX GPUs.
@@ 4.2 Operator Adaptation
The efficient adaptation of SpikingBrain large models to MetaX GPUs relies on the comprehensive MetaX software ecosystem. The overall process can be divided into two major components: Triton adaptation and CUDA migration to the MetaX self-developed MACA framework. These two pathways target different operator subsets in SpikingBrain models. While distinct in implementation, they complement each other and together constitute a complete hardware adaptation framework tailored for MetaX GPUs (see Figure 5).
@@
  • JIT Compilation Optimization: During Triton’s just-in-time compilation stage, the linear attention operators (Yang et al., 2024c) used in SpikingBrain are reorganized at the code level. By refining instruction pipelining and register allocation, the operators achieve a balance between memory access latency and computational density. This stage directly leverages Triton’s hardware-agnostic general optimization strategies, enabling dynamic and efficient execution of compute kernels.
@@
  • Target Code Generation via MetaX Compiler: Finally, Triton kernels are transformed by the MetaX compiler backend into executable target machine code for MetaX GPUs. Beyond standard code generation, the compiler introduces deep optimizations for tensor cores, SIMD instruction sets, and memory alignment constraints, ensuring operators run at near-peak hardware performance.
@@
It is worth emphasizing that the entire adaptation and migration process for brain-inspired large models on MetaX GPUs adheres to a user-friendly design philosophy. Whether through Triton kernel optimization or CUDA-to-MACA migration, end users can preserve their existing programming practices and interface calls. In most cases, no extensive code modifications are required for the models to run efficiently on MetaX GPUs. Meanwhile, the MetaX ecosystem provides unified debugging and performance analysis tools, enabling developers to transparently monitor execution characteristics on hardware and tune performance as needed.
@@ 4.3 Parallel Topology
Data parallelism (Valiant, 1990) involves partitioning the training data into batches, each processed by a separate GPU. Each GPU maintains a complete replica of the model and performs forward and backward passes independently, while gradient synchronization occurs only at the beginning of the backward propagation. Data parallelism offers low communication overhead and excellent scalability. As computing resources increase, the scale of data parallelism can grow accordingly, increasing the total batch size to enhance overall throughput and reduce training time. During training, we employ ZeRO (Rajbhandari et al., 2020) to eliminate redundancy of the optimizer states, effectively distributing the GPU memory pressure.
@@
Pipeline parallelism divides the model into different stages by layers. Each GPU is assigned a subset of layers, and intermediate activations are communicated between GPUs during the forward and backward passes using peer-to-peer (p2p) communication. We employ the 1F1B (Narayanan et al., 2021; Harlap et al., 2018) scheduling algorithm for efficient pipeline execution. Furthermore, incorporating Mixture-of-Experts (MoE) layers interspersed with dense FFN layers in SpikingBrain-76B helps mitigate the memory imbalance issue, where the first pipeline stage typically exhibits higher memory usage compared to other stages.
@@
In Mixture-of-Experts models, most parameters reside within the expert networks. Expert parallelism (Rajbhandari et al., 2022) addresses this by partitioning experts across multiple GPUs. This requires all-to-all communication to dispatch and combine tokens. During the training of SpikingBrain-76B, we employ Grouped GEMM kernels to accelerate computation across experts, and an auxiliary loss is used to promote balanced token routing across experts.
@@
Training with long sequences incurs substantial memory costs for storing intermediate activations. Sequence parallelism tackles this by splitting the input sequence along its length dimension, with each GPU processes a contiguous chunk. Communication primitives like all-to-all (a2a) or peer-to-peer (p2p) are then used to exchange necessary information between devices, enabling the computation of attention over the entire sequence. For our experiments, we employ DeepSpeed Ulysses (Jacobs et al., 2023) for efficient sequence parallelism and incorporate specialized algorithms for linear attention modules to further improve performance. For linear attention, AllGather communication is used to collect intermediate local states (Sun et al., 2025) when the SP size is small, while for larger SP sizes (e.g., across multiple nodes), the All-Scan primitive from ZeCO (Chou et al., 2025) sequence parallelism is adopted to effectively eliminate communication overhead.
@@
The SpikingBrain-7B model is trained using the Colossal-AI (Li et al., 2023b) framework. For 128k sequence length training, a parallel strategy of 32-way data parallelism and 8-way sequence parallelism is adopted. To reduce GPU memory consumption, techniques such as ZeRO-2 (Rajbhandari et al., 2020) and activation recomputation (Chen et al., 2016; Korthikanti et al., 2022) are applied. Sequence parallelism is implemented with DeepSpeed Ulysses (Jacobs et al., 2023), which performs all-to-all communication within nodes. In addition, the number of attention heads is padded to satisfy the partition requirements.
The SpikingBrain-76B model is trained using the Megatron (Shoeybi et al., 2019) framework. For 8k sequence length training, the parallel configuration consists of 128-way data parallelism, 8-way expert parallelism, and 4-way pipeline parallelism. ZeRO (Rajbhandari et al., 2020) and selective activation recomputation (Chen et al., 2016; Korthikanti et al., 2022) are applied to reduce memory consumption. For extended contexts of 32k and 128k, 4-way and 8-way sequence parallelism are added on top of the base 8k parallel strategy, respectively. The 76B hybrid model integrates linear attention and (local) softmax attention. The softmax attention branch uses DeepSpeed Ulysses (Jacobs et al., 2023) with all-to-all communication and head padding to enable partitioning, while the linear attention branch employs the AllGather-based SP algorithm (LASP-2 (Sun et al., 2025)) to reduce communication overhead.
@@ 5.1 Downstream Performance
We conduct a comprehensive evaluation of the two converted efficient models, SpikingBrain-7B and -76B, across a broad set of downstream tasks [^3] This section evaluates the effectiveness of the architectural design and conversion pipeline. All models remain in floating-point format; results for spiking models are presented in Section 5.5.. They are compared against open-source baselines of similar scale, including Hybrid, Transformer, and MoE architectures, as well as the strong base model Qwen2.5-7B. To ensure fairness, all evaluations are performed under identical settings using the OpenCompass (Buitrago & Nystrom, 2019) framework. In selecting evaluation metrics, we place greater emphasis on pretraining-oriented benchmarks (e.g., MMLU, C-Eval), as these better indicate whether our models—trained with fewer than 200B tokens during efficient conversion—can inherit the knowledge and modeling capacity of the base model. Moreover, these benchmarks are less confounded by data quality factors such as alignment performance after SFT. Note that, except for Qwen2.5, the other baselines are trained on limited Chinese data, resulting in disadvantages on CMMLU and C-Eval. For more information on the baselines and benchmarks, please refer to Appendix A.
Pre-trained Models
As shown in Table 1, our SpikingBrain-7B linear model recovers nearly 90% of the base model’s performance across benchmarks, reaching a level comparable to advanced Transformer models such as Mistral-7B and Llama3-8B. This indicates that, with appropriate training strategies, linear attention architectures can preserve strong general-purpose modeling ability while substantially reducing complexity. However, due to the significant structural modifications of the attention mechanism, there remains a certain performance gap between the 7B pure linear model and base Qwen2.5-7B. This suggests that the extent of recovery is strongly influenced by architectural changes when pursuing extreme efficiency.
Table 1: Performance evaluation of the SpikingBrain-7B pre-trained model. All models are tested with the HuggingFace framework and evaluated using a perplexity-based method.
@@
MMLU 65.84 63.24 62.56 58.19 65.74 74.21
CMMLU 71.58 42.50 44.58 38.42 52.44 81.73
ARC-C 43.32 47.53 45.13 37.18 51.96 44.04
HS 70.89 71.50 75.81 52.02 71.60 72.81
Ceval 69.80 41.93 47.04 36.40 51.46 81.60
@@
Table 2: Performance evaluation of the SpikingBrain-76B pre-trained model. All models are tested with the vLLM (Kwon et al., 2023) framework and evaluated using a perplexity-based method.
@@
MMLU 73.58 67.17 71.23 69.57 75.94 75.31
CMMLU 78.83 51.11 52.70 52.94 61.80 81.50
ARC-C 42.00 49.10 48.98 46.21 57.64 43.56
HS 73.31 79.39 79.42 79.15 79.34 73.37
Ceval 78.89 48.94 54.39 49.26 61.02 81.68
Instruct Models
@@
Long-context Capability
To assess real-world long-context modeling and understanding, we evaluate SpikingBrain-76B (after stage‑2 SFT) on six LongBench tasks (Bai et al., 2024). Given the baseline’s maximum positional embeddings, we truncate prompts to 32k. As shown in Table 5, our model is competitive with other hybrid and Transformer baselines while using only four global softmax attention layers.
5.2 Long-context Efficiency
Beyond task performance, we also place particular emphasis on the efficiency of our models in long-sequence inference. Benefiting from the integration of (hybrid) linear attention with the MoE mechanism, both SpikingBrain models exhibit substantial inference speedups. Since the 7B model is more compact and shows stronger efficiency gains, we adapt it to the HuggingFace framework with multi-device sequence-parallel inference (ZeCO + p2p), successfully supporting efficient prefill up to 4M tokens. As shown in Figure 7 and Table 6, at an input length of 1M tokens, the SpikingBrain-7B model achieves a 26.5× speedup in TTFT (Time to First Token) compared with the Qwen2.5 baseline using full attention and all-to-all communication. Due to resource limitations and the number of attention heads, Qwen2.5 is difficult to evaluate at 4M tokens. Based on the fitted scaling curve, we conservatively estimate a speedup of over 100×. Moreover, as sequence length and GPU count increase proportionally, the time overhead of our 7B model remains nearly constant. These results demonstrate the scalability and superiority of our approach in ultra-long-sequence scenarios, providing strong support for the deployment of more efficient large language models in practical applications.
For long-sequence training, we further measure Throughput per GPU Second (TGS) on SpikingBrain-7B and the Qwen2.5 baseline under the sequence parallelism framework. The experiments show a 5.36× throughput advantage for our 7B model at a sequence length of 128k, consistent with the TTFT improvements observed during inference and attributable to its efficient attention design.
5.3 CPU-side Inference
In this section, we deploy compressed 1B-parameter models, including the SpikingBrain linear model with SWA and GLA, as well as Llama3.2 (Dubey et al., 2024), on a CPU-based mobile inference framework. The deployment leverages llama.cpp as the inference backend. As illustrated in Figure 8, the workflow consists of four main steps: i) Weight conversion and quantization, ii) Model registration and tensor mapping, iii) Graph and operator optimization, and iv) Quantized inference. See Appendix A.6 for details.
5.4 Performance on the MetaX Cluster
The performance of the MetaX C550 GPU cluster is evaluated from two perspectives: training efficiency and system stability. For the SpikingBrain-7B model during the training phrase, we achieve a TGS (Token Per Second) of 1558 and an MFU (Model FLOPs Utilization) of 23.4% (8-way DP, 4-way PP, PP micro-batch size 2, global batch size 512), reflecting high computational efficiency and effective resource utilization.
Overall, these results highlight the capability of the MetaX GPU cluster to efficiently and reliably support large-scale long-duration training tasks, enabling stable training of models with tens of billions of parameters.
5.5 Analysis of Spiking Scheme
Scheme Deployment and Accuracy Maintenance
Winogrande Arc-e
Arc-c
HellaSwag
PIQA MMLU CMMLU Avg.
SpikingBrain-7B -
W8ASpike
0.6895 0.7861 0.5410 0.6758 0.7979 0.6546 0.6677 0.6875
SpikingBrain -76B-
W8ASpike
0.7148 0.7949 0.5371 0.6863 0.8004 0.7081 0.7512 0.7133
Spike Distribution Characteristics
To evaluate the sparsity and efficiency of the proposed spiking scheme, we analyze the statistical firing patterns under the bitwise–ternary coding method. We collect the absolute values of input spike integers across all linear projection layers. The detailed distribution of spike counts is illustrated in Figure 9. Take the SpikingBrain-7B model for example, results show that 94.1% of values fall within [0,7], and only about 1% exceed 16. After expansion into bitwise–ternary spikes, the average number of spikes fired per channel is only 1.13, while 18.4% of channels do not fire any spikes at all. Since 94.1% of values are within 7 (i.e., within the INT3 range), we further analyze the firing rate under a step size of 3. By filling inactive bits within the 3-step window with 0, the overall sparsity reaches about 69.15%.
Spike Visualization Interface
Energy Efficiency with Asynchronous Hardware
For computation, we estimate and compare the energy consumption of MAC (multiply–accumulate) operations under different paradigms, based on published hardware energy consumption data at 45nm technology (Yao et al., 2024b):
    $E=\text{Average Spikes}\times E_{\text{INT8 Add}}.$ (24)
6 Conclusion
This work provides a comprehensive demonstration of efficient brain-inspired large model training and inference on the MetaX GPU cluster. By integrating several key techniques—including novel (hybrid) linear architectures beyond Transformers, sparse MoE design, adaptive-threshold sparse spiking activation, and lightweight conversion training—we validate and implement a practical development pipeline for spiking-based large models on a non-NVIDIA cluster with hundreds of GPUs.
Author Contributions
Yuqi Pan and Yuhong Chou designed the model architectures. Siyu Ding proposed the adaptive threshold spiking neuron and conducted spiking-based training. Han Xu and Xuerui Qiu developed three spike coding techniques. Yuqi Pan, Yupeng Feng, and Zehao Liu carried out the framework adaptation on MetaX and model training. Jinghao Zhuang, Yupeng Feng, Zehao Liu, Yuqi Pan, Bohan Sun, and Yuhong Chou performed model evaluation. Anjie Hu and Anlin Deng contributed to code adaptation and refinement. Peng Zhou performed CPU adaptation and model testing. Yuqi Pan, Shurong Wang, Guoqi Li, Man Yao, Jibin Wu, Jian Yang, Guoliang Sun, and Bo Xu contributed to the discussion and analysis of experimental results. The project was supervised by Bo Xu and Guoqi Li.
Acknowledgements
Appendix A Experiments
A.1 Benchmarks
In selecting evaluation metrics, we place greater emphasis on pretraining-oriented general-purpose benchmarks: MMLU (Hendrycks et al., 2020), CMMLU (Li et al., 2023a), C-Eval (Huang et al., 2023), ARC-C (Clark et al., 2018), and HS (Zellers et al., 2019), as these better indicate whether our models—trained with fewer than 200B tokens during efficient conversion—can inherit the knowledge and modeling capacity of the base model. Moreover, these benchmarks reveal whether the converted models preserve strong generalization while demonstrating significant long-sequence efficiency advantages, without being confounded by data quality factors such as alignment performance after SFT.
For SpikingBrain chat models after three-stage SFT, we evaluate general knowledge, long-sequence modeling, and instruction following, additionally including NQ (Kwiatkowski et al., 2019), TrQ (Joshi et al., 2017), QuALITY (Pang et al., 2022), and IFEval (Zhou et al., 2023). To assess real-world long-context modeling and understanding capability, we also evaluate six popular tasks on LongBench (Bai et al., 2024).
A.2 Models
We compare SpikingBrain models against open-source baselines of comparable scale spanning Hybrid, Transformer, and MoE architectures:
  • Linear-complexity models: Falcon-Mamba-7B (Zuo et al., 2024), Mistral-7B (team, 2023).
  • Hybrid linear-complexity models: Zamba-v1-7B (Glorioso et al., 2024), Jamba-52B-A12B (Lieber et al., 2024).
  • Quadratic-complexity models: Llama3.1-8B (Dubey et al., 2024), Qwen2.5-7B (Yang et al., 2024b), Mixtral-47B-A13B (Jiang et al., 2024), LLama2-70b (Touvron et al., 2023), Gemma2-27B (Team et al., 2024).
A.3 Evaluation of Instruct Models
We apply three-stage SFT to the converted models to add instruction following. As shown in Table 4, the aligned models match similarly sized open-source chat baselines across general knowledge, long-sequence modeling, and instruction following, while avoiding overfitting and preserving pretrained capabilities—indicating that (hybrid) linear architectures remain stable and scalable under alignment.